# SpecVLM: Fast Speculative Decoding in Vision-Language Models

### v1

@@ Abstract
Speculative decoding is a powerful way to accelerate autoregressive large language models (LLMs), but directly porting it to vision-language models (VLMs) faces unique systems constraints: the prefill stage is dominated by visual tokens whose count scales with image resolution and video length, inflating both compute and memory—especially the key-value (KV) cache. We study speculative decoding for VLMs and introduce SpecVLM, a practical system that (1) establishes a strong EAGLE-2-style baseline, EagleVLM, delivering 1.5–2.3 end-to-end speedups over full autoregressive inference, and (2) further accelerates VLM inference with an elastic visual compressor that adaptively selects among pruning, pooling, convolution, and resampler primitives to balance FLOPs/parameters and accuracy per input. To avoid costly offline distillation corpora, we propose an online-logit distillation protocol that trains the draft model with on-the-fly teacher logits and penultimate features using a combined cross-entropy and Smooth L1 objective, eliminating storage and preprocessing while remaining compute-efficient. This protocol reveals a training-time scaling effect: longer online training monotonically increases the draft model’s average accepted length, improving speculative efficiency. Empirically, SpecVLM achieves additional acceleration, culminating in 2.5–2.9 end-to-end speedups within 5 epochs across LLaVA and MMMU, consistently over resolutions and task difficulties, while preserving the target model’s output distribution (lossless decoding). Our code is available at https://github.com/haiduo/SpecVLM.
@@ 1 Introduction
To tackle these dual bottlenecks, we combine two complementary accelerations. First, we apply speculative decoding (Leviathan et al., 2023) to reduce the number of autoregressive steps, improving latency while preserving the target model’s output distribution (lossless decoding). Second, we integrate visual token compression (Li et al., 2024a) and propose an elastic, question-adaptive compressor to reduce the number of visual tokens, which alleviates both compute and memory in the draft model prefill and also reduces KV-cache traffic during decoding. A comparison is shown in Figure 1(b).
@@
Speculative decoding—running a smaller, faster draft model to propose multiple tokens and verifying them with a heavier target model—has proven effective for text-only LLMs (e.g., EAGLE-style (Li et al., 2024c; d) methods). However, porting this to VLMs introduces unique challenges (see Figure 2). First, the draft must process visual inputs efficiently during prefill; naively reusing LLM drafts leaves the draft burdened by large visual token sets. Second, our survey shows four major classes of visual token compressors—pruning (Shang et al., 2024; Chen et al., 2024; Huang et al., 2024; Zhang et al., 2024; Cao et al., 2024; Alvar et al., 2025; Chen et al., 2025; Tao et al., 2025; Ye et al., 2025a; Lin et al., 2025; Ye et al., 2025b), pooling (Cha et al., 2024; Li et al., 2024b; Cai et al., 2024a), convolutional (Chu et al., 2023; Dong et al., 2024), and resampler-based (Li et al., 2023; Fan et al., 2024; Hu et al., 2024; Li et al., 2024a; 2025a; Zhang et al., 2025)—each with complementary strengths and weaknesses. Parameter-free methods (pruning, pooling) are cheap but less expressive; learned resamplers and convolutional compressors are more expressive but add parameters and compute. There has been little exploration of how to select these compressors inside a speculative decoding pipeline to match task difficulty, model capacity, and compute budgets.
@@
In this work, we take a comprehensive approach. We first construct EagleVLM, a strong EAGLE-2-inspired speculative baseline for VLMs that integrates a faithful draft with the full target model, demonstrating 1.5–2 speedups over end-to-end autoregressive inference. Building on this baseline, we introduce two key innovations to further reduce VLM inference cost: (1) an elastic visual compressor that dynamically selects pruning, pooling, resampler, and convolution operators to meet accuracy-compute trade-offs per input; and (2) an online-logit distillation protocol that trains the draft with on-the-fly teacher supervision, avoiding large offline distillation corpora while improving alignment.
@@
An empirical finding is particularly noteworthy: with fixed training data and draft architecture, increasing online training time monotonically increases the draft model’s average accepted length, thereby improving speculative efficiency. This training-time scaling effect suggests multimodal speculative decoding benefits from longer, targeted training regimes beyond what LLM-only studies have reported. Finally, our SpecVLM—combining the elastic compressor with online-logit training—delivers overall end-to-end acceleration up to 2.5–2.9 within 5 epochs, without sacrificing output fidelity.
@@ 2.1 Speculative Decoding
Speculative decoding accelerates autoregressive models by letting a smaller, faster draft propose multiple tokens that a larger target verifies in parallel (Leviathan et al., 2023; Chen et al., 2023). Early work explored greedy blockwise acceptance (Stern et al., 2018; Sun et al., 2021), and subsequent methods generalized to non-greedy settings (Chen et al., 2023). Systems such as SpecInfer (Miao et al., 2023), Medusa (Cai et al., 2024b), and Hydra (Ankner et al., 2024) improve efficiency via parallel verification and auxiliary heads, but can suffer from limited token diversity and coupling across steps. EAGLE (Li et al., 2024c; d) addresses this by autoregressing at the feature level and employing static/dynamic tree-structured drafts that decouple time steps. Jakiro (Huang et al., 2025a) further enhances diversity with Mixture-of-Experts (MoE) heads at the cost of added latency. However, these approaches largely assume text-only models and overlook VLM-specific constraints: heavy visual prefill, large KV caches, and resolution-/video-driven sequence growth. They also typically depend on offline teacher-logit corpora, which are cumbersome at multimodal scales. In contrast, SpecVLM uses a single-layer Transformer decoder as the draft and trains it with online-logit distillation, avoiding offline data generation while preserving lossless decoding.
@@ 3.1 Preliminaries: Bottleneck Analysis
Here, we define as the average accepted tokens per SD round. Under a simplifying independence assumption with a per-token acceptance probability , a common approximation is . The denominator in Eq. equation 2 has three terms: (i) the relative latency of the draft model (), (ii) the cost of multi-token verification (), and (iii) the negligible cost of speculative sampling (). In practice, speculative decoding tends to be memory-bound (Sadhukhan et al., 2024) at small batch sizes because multi-token verification can saturate memory bandwidth before compute is fully utilized. Hence (ii) and (iii) are usually small; the speedup is most sensitive to the draft latency and the average accepted length .
@@ 3.4 Elastic Vision Compressor
(c) Dynamic selection of compression experts. Evaluating all branches every time is sub-optimal on resource-limited devices. We therefore introduce a question-aware selector that chooses one operator per input based on the gate’s top-1 logit, considering question difficulty and image-text relevance. We also include a Text-only branch (drop all vision tokens) to exploit LLMs’ zero-shot ability when the question can be answered without visual evidence. This dynamic selection is the default strategy in SpecVLM and yields favorable accuracy-compute trade-offs. Since selection is discrete and non-differentiable, we use Gumbel-Softmax (Jang et al., 2016) to approximate the categorical distribution during training; at inference, sampling is replaced by without extra overhead.
@@ 4.1 Experiment Setup
Training and Testing Setup. We freeze the target VLM and train only the draft model’s decoder layer and the vision compressor (Figure 3b). For LLaVA-1.5-7B/13B, we adopt the 655K multimodal instruction-following dataset from official LLaVA-1.5 visual instruction tuning (150K GPT-generated instructions + 515K academic VQA). For LLaVA-1.6, since the original dataset is unavailable, we use the Open-LLaVA-NeXT construction (1.02M samples). We train for 1 epoch with AdamW (), batch size 128, no warmup; learning rate is 8e-5 for 7B and 5e-5 for 13B (weight decay 0). Training runs on 8 AMD MI250 GPUs; one epoch takes 1̃5/20 hours for LLaVA-1.5-7B/13B and 2̃6/35 hours for LLaVA-1.6-7B/13B. We test on AMD Instinct MI250-64G and NVIDIA A100-80G. All latency is measured on a single GPU to ensure consistency across target sizes. Following EAGLE-2, the draft uses tree attention with Top-K=10, total tokens=60, depth=7.
@@ 5.1 Component Analysis
We ablate the contribution of EagleVLM, the elastic visual compressor, and online-logit distillation on LLaVA-1.6-7B (Table 3). Configurations include: (1) EagleVLM (EAGLE-2-style speculative baseline); (2) Weight (weighted fusion of compression experts); (3) Concat (multi-granularity feature concatenation); and (4) Select (dynamic expert selection; default SpecVLM).
The results show cumulative gains. Starting from EagleVLM, adding weighted fusion improves both and . Multi-granularity concatenation preserves these gains. Dynamic selection (full SpecVLM) yields the best overall performance, reaching 2.11 average speedup and 3.67 average accepted length. These findings highlight the importance of adaptability—particularly the question-aware expert selection—in achieving robust speed-accuracy trade-offs.
@@
Analysis of Isolated Branches. Pruning and pooling, which are parameter-free, maintain robustness up to a 10-20 compression ratio, improving while keeping high. The resampler performs best with 2 queries, as more queries add overhead without boosting . Lightweight convolutions match pooling at 3–5 ratios, but both and drop at ratios of 10 or more. These insights guide our strategy: use 10-20 pooling/pruning for simple or text-based tasks, 3 convolution and a 2-query resampler for detailed visual information inputs.
@@ 5.2 Training-time Scaling
We analyze how extended online training impacts speculative efficiency. As shown in Figures 5 and 6, both accepted length () and speedup () increase monotonically with epochs across all variants and sizes. This indicates that longer online training improves draft-target alignment, increases accepted tokens per round, and amplifies acceleration.
Notably, these results underscore the potential of online-logit distillation for VLM speculative decoding: substantial efficiency gains arise solely from longer training, without architectural changes or larger draft capacity. The effect is consistent across sizes, supporting scalability.
By epoch 5, the draft had not converged. Extending training would significantly increase compute; given limited resources, we defer further epochs. We hypothesize convergence by around the 8th epoch and plan to update results when resources allow.

### v2

@@ Abstract
Speculative decoding is a powerful way to accelerate autoregressive large language models (LLMs), but directly porting it to vision-language models (VLMs) faces unique systems constraints: the prefill stage is dominated by visual tokens whose count scales with image resolution and video length, inflating both compute and memory—especially the key-value (KV) cache. We study speculative decoding for VLMs and introduce SpecVLM, a practical system that (1) establishes a strong EAGLE-2-style baseline, EagleVLM, delivering 1.5–2.3 end-to-end speedups over full autoregressive inference, and (2) further accelerates VLM inference with an elastic visual compressor that adaptively selects among pruning, pooling, convolution, and resampler primitives to balance FLOPs/parameters and accuracy per input. To avoid costly offline distillation corpora, we propose an online-logit distillation protocol that trains the draft model with on-the-fly teacher logits and penultimate features using a combined cross-entropy and Smooth L1 objective, eliminating storage and preprocessing while remaining compute-efficient. This protocol reveals a training-time scaling effect: longer online training monotonically increases the draft model’s average accepted length, improving speculative efficiency. Empirically, SpecVLM achieves additional acceleration, culminating in 2.5–2.9 end-to-end speedups within 5 epochs across LLaVA and MMMU, consistently over resolutions and task difficulties, while preserving the target model’s output distribution (lossless decoding).
@@ 1 Introduction
To tackle these dual bottlenecks, we combine two complementary accelerations. First, we apply speculative decoding (Leviathan et al., 2023) to reduce the number of autoregressive steps, improving latency while preserving the target model’s output distribution (lossless decoding). Second, we integrate visual token compression (Li et al., 2024a) and propose an elastic, question-adaptive compressor to reduce the number of visual tokens. This alleviates both compute and memory usage in the draft model prefill and reduces KV-cache traffic during decoding, as shown in Figure 1(b).
@@
Speculative decoding—running a smaller, faster draft model to propose multiple tokens and verifying them with a heavier target model—has proven effective for text-only LLMs. However, porting this to VLMs introduces unique challenges (see Figure 2). First, the draft must process visual inputs efficiently during prefill; naively reusing LLM drafts leaves the draft burdened by large visual token sets. Second, our survey shows four major classes of visual token compressors—pruning (Shang et al., 2024; Chen et al., 2024; Huang et al., 2024; Zhang et al., 2024; Cao et al., 2024; Alvar et al., 2025; Chen et al., 2025; Tao et al., 2025; Ye et al., 2025a; Lin et al., 2025; Ye et al., 2025b), pooling (Cha et al., 2024; Li et al., 2024b; Cai et al., 2024a), convolutional (Chu et al., 2023; Dong et al., 2024), and resampler-based (Li et al., 2023; Fan et al., 2024; Hu et al., 2024; Li et al., 2024a; 2025a; Zhang et al., 2025)—each with complementary strengths and weaknesses. Parameter-free methods (pruning, pooling) are cheap but less expressive; learned resamplers and convolutional compressors are more expressive but add parameters and compute. There has been little exploration of how to select or combine these compressors inside a speculative decoding pipeline to match task difficulty, model capacity, and compute budgets.
@@
We take a comprehensive approach. We first construct EagleVLM, a strong EAGLE-2-inspired speculative baseline for VLMs that integrates a faithful draft with the full target model, demonstrating 1.5–2 speedups over end-to-end autoregressive inference. Building on this baseline, we introduce two key innovations to further reduce VLM inference cost: (1) an elastic visual compressor that dynamically selects pruning, pooling, resampler, and convolution operators to meet accuracy-compute trade-offs per input; and (2) an online-logit distillation protocol that trains the draft with on-the-fly teacher supervision, avoiding large offline distillation corpora while improving alignment.
@@
An empirical finding is particularly noteworthy: with fixed training data and draft architecture, increasing online training time monotonically increases the draft model’s average accepted length, thereby improving speculative efficiency. This training-time scaling effect suggests multimodal speculative decoding benefits from longer, targeted training regimes beyond what LLM-only studies have reported. Finally, our SpecVLM—combining the elastic compressor with online-logit training—delivers overall end-to-end acceleration up to 2.5–2.9 within 5 epochs, without sacrificing output fidelity. Our contributions are summarized as follows:
@@ 2.1 Speculative Decoding
Speculative decoding accelerates autoregressive models by letting a smaller, faster draft propose multiple tokens that a larger target verifies in parallel (Leviathan et al., 2023; Chen et al., 2023). Early work explored greedy blockwise acceptance (Stern et al., 2018). Systems such as SpecInfer (Miao et al., 2023), Medusa (Cai et al., 2024b), and Hydra (Ankner et al., 2024) improve efficiency via parallel verification and auxiliary heads, but can suffer from limited token diversity and coupling across steps. EAGLE (Li et al., 2024c; d) addresses this by autoregressing at the feature level and employing static/dynamic tree-structured drafts that decouple time steps. Jakiro (Huang et al., 2025a) further enhances diversity with Mixture-of-Experts (MoE) heads at the cost of added latency. However, these approaches largely assume text-only models and overlook VLM-specific constraints: heavy visual prefill, large KV caches, and resolution-/video-driven sequence growth. They also typically depend on offline teacher-logit corpora, which are cumbersome at multimodal scales. Recent works have explored speculative decoding in VLMs through SPD-MLLM (Gagrani et al., 2024) and Spec-LLaVA (Huo et al., ), but these rely on carefully designed standalone LLM drafts requiring significant architectural modifications. TwigVLM (Shao et al., 2025) combines speculative decoding with pruning techniques but produces lossy results. Additionally, mainstream speculative decoding approaches rely on offline dataset generation (Li et al., 2024c; d; Cai et al., 2024b; Ankner et al., 2024; Huang et al., 2025a), which can be storage-intensive and computationally expensive for large VLMs. In contrast, our SpecVLM uses only a single transformer decoder layer combined with online logits distillation, eliminating the cumbersome data preprocessing.
@@ 3.1 Preliminaries: Bottleneck Analysis
Here, we define as the average accepted tokens per SD round. Under a simplifying independence assumption with a per-token acceptance probability , a common approximation is . The denominator in equation 2 has three terms: (i) the relative latency of the draft model (), (ii) the cost of multi-token verification (), and (iii) the negligible cost of speculative sampling (). In practice, speculative decoding tends to be memory-bound (Sadhukhan et al., 2024) at small batch sizes because multi-token verification can saturate memory bandwidth before compute is fully utilized. Hence (ii) and (iii) are usually small; the speedup is most sensitive to the draft latency and the average accepted length .
@@ 3.4 Elastic Vision Compressor
(c) Dynamic selection of compression experts. Evaluating all branches every time is sub-optimal on resource-limited devices. We introduce a question-aware selector that chooses one operator per input based on the gate’s top-1 logit, considering question difficulty and image-text relevance. We also include a Text-only branch (drop all vision tokens) to exploit LLMs’ zero-shot ability when the question can be answered without visual evidence. This dynamic selection is the default strategy in SpecVLM and yields favorable accuracy-compute trade-offs. Since selection is discrete and non-differentiable, we use Gumbel-Softmax (Jang et al., 2016) to approximate the categorical distribution during training; at inference, sampling is replaced by without extra overhead.
@@ 4.1 Experiment Setup
Training and Testing Setup. We freeze the target VLM and train only the draft model’s decoder layer and the vision compressor (Figure 3b). For LLaVA-1.5-7B/13B, we adopt the 655K multimodal instruction-following dataset from official LLaVA-1.5 visual instruction tuning (150K GPT-generated instructions + 515K academic VQA). For LLaVA-1.6, since the original dataset is unavailable, we use the Open-LLaVA-NeXT construction (1.02M samples). We train for 1 epoch with AdamW (), batch size 128, no warmup; learning rate is 8e-5 for 7B and 5e-5 for 13B (weight decay 0). Training runs on 8 AMD MI250 GPUs; one epoch takes 15/20 hours for LLaVA-1.5-7B/13B and 26/35 hours for LLaVA-1.6-7B/13B. We test on AMD Instinct MI250-64G and NVIDIA A100-80G. All latency is measured on a single GPU to ensure consistency across target sizes. Following EAGLE-2, the draft uses tree attention with Top-K=10, total tokens=60, depth=7.
@@ 5.1 Component Analysis
We ablate the contribution of EagleVLM, the elastic visual compressor, and online-logit distillation on LLaVA-1.6-7B (Table 3). Configurations include: (1) EagleVLM (EAGLE-2-style speculative baseline); (2) Weight (weighted fusion of compression experts); (3) Concat (multi-granularity feature concatenation); and (4) Select (dynamic expert selection; default SpecVLM). The results show cumulative gains. Starting from EagleVLM, adding weighted fusion improves both and . Multi-granularity concatenation preserves these gains. Dynamic selection (full SpecVLM) yields the best overall performance, reaching 2.03 average speedup and 3.69 average accepted length. These findings highlight the importance of adaptability—particularly the question-aware expert selection—in achieving robust speed-accuracy trade-offs.
@@
| Prun | 30 | 2.00 | 3.36 | Pool | 30 | 2.01 | 3.37 | Conv | 30 | 1.67 | 2.73 | Resamp | 20 | 1.99 | 3.36 |
Analysis of Isolated Branches. Pruning and pooling, which are parameter-free, maintain robustness up to a 10-20 compression ratio, improving while keeping high. The resampler performs best with 2 queries, as more queries add overhead without boosting . Lightweight convolutions match pooling at 3-5 ratios, but both and drop at ratios of 10 or more. These insights guide our strategy: use 10-20 pooling/pruning for simple or text-based tasks, 3 convolution and a 2-query resampler for detailed visual information inputs.
@@ 5.2 Training-time Scaling
Figures 5 and 6 illustrate that both accepted length () and speedup () consistently rise with each epoch across all variants and sizes. This suggests that prolonged online training enhances draft-target alignment, boosts accepted tokens per round, and increases acceleration. Notably, these results underscore the potential of online-logit distillation in improving VLM speculative decoding, without requiring architectural changes or increased draft capacity. This effect is consistently observed across various model sizes, demonstrating strong scalability. By the 5th epoch, the draft model starts to converge, achieving optimal performance by the 8th epoch. However, further extending the training period would substantially raise computational costs, so we selected the 5th epoch as the optimal trade-off point.